OpenStudy (anonymous):

turn 10sqrt(10)+11sqrt(11) into the following exact answer: sqrt( 2331+ 220sqrt(110))

OpenStudy (anonymous):

\[10\sqrt{10}+11\sqrt{11} = \sqrt{2331+220\sqrt{110}}\]